OpenStudy (winsteria):

MEDAL AND FAN A porch has the shape of a quarter circle with a diameter of 5 ft. What is the approximate area of the porch, rounded to the nearest square foot? Use 3.14 to approximate pi star.

OpenStudy (winsteria):

@sammixboo @iGreen @Legends

syed98 (syedmohammed98):

the formula for quarter circle is 1/4 * 3.14 * r so it's 1/4 * 3.14 * (2.5)

OpenStudy (cggurumanjunath):

( pi / 4 *4 ) d*d

OpenStudy (winsteria):

1.9625?

OpenStudy (cggurumanjunath):

( 3 . 14 / 16 ) * 5 * 5

syed98 (syedmohammed98):

do u have options in the question?

OpenStudy (winsteria):

no

OpenStudy (cggurumanjunath):

4.90625 foot ^ 2 ==== = Final Answer !

syed98 (syedmohammed98):

http://prntscr.com/6nclik

OpenStudy (cggurumanjunath):

Area of the porch !

OpenStudy (winsteria):

I don't know :/

syed98 (syedmohammed98):

oh shhhhhhhit.....sorryyyyyy.......hey CGG IS RIGHT ......i thought it's pouch.....lol really sorry

OpenStudy (cggurumanjunath):

no dought ! 4.90625 foot ^2 is perfect answer !

syed98 (syedmohammed98):

agreed @CGGURUMANJUNATH

OpenStudy (winsteria):

9.8125?

syed98 (syedmohammed98):

CGG gave the answer :)

OpenStudy (winsteria):

oh so it's 4.90625?

syed98 (syedmohammed98):

yup =)

OpenStudy (winsteria):

Thank You

syed98 (syedmohammed98):

ur welcome

OpenStudy (winsteria):

It can't be the big answer though...

OpenStudy (winsteria):

4.9?

OpenStudy (winsteria):

@StudyGurl14

OpenStudy (winsteria):

@AlexandervonHumboldt2 @Nurali

OpenStudy (winsteria):

@ribhu

OpenStudy (igreen):

\(A = \pi r^2\) We are given the diameter, 5. Radius is half the diameter. 5 / 2 = ?

OpenStudy (winsteria):

2.5

OpenStudy (igreen):

Yes, so we have: \(A = (3.14)(2.5^2)\) Simplify

OpenStudy (winsteria):

19.625?

OpenStudy (igreen):

Yes, now divide by 4.

OpenStudy (winsteria):

4.90625

OpenStudy (igreen):

Yes, now round it to the nearest whole number.

OpenStudy (winsteria):

49?

OpenStudy (igreen):

No, 4.90625 rounds to 5.

OpenStudy (winsteria):

Thank You

jabez177 (jabez177):

@gabbyalicorn I just had to google it. xD

OpenStudy (gabbyalicorn):

:/ i did 2

OpenStudy (gabbyalicorn):

wait so its 5? :3

jabez177 (jabez177):

So yeah. 5 is the right answer. I DID THIS LESSON!

OpenStudy (gabbyalicorn):

just 5 hmmm kay.

jabez177 (jabez177):

I mastered the checkpoint. I just saw...